OpenStudy (anonymous):

Solve the system of equations x - y = 10 2x + y = 2

OpenStudy (anonymous):

Solving Simultaneously: X=-4 & Y=6.

OpenStudy (nubeer):

re- arrange your first equation x = 10+y (got by adding y on both sides.) now put this value of x in the second equation

OpenStudy (jiteshmeghwal9):

x-y=10 x=10+y Now putting the value in the next equation 2(10+y)+y=2 20+2y+y=2 20+3y=2 3y=-18 y=-6 Now put the value of y in eq.(x=10+y) to get x :)

OpenStudy (zpupster):

they used substitution method which is fine but you can add these equations since -y and y = 0 giving you x only. x - y = 10 2x + y = 2 ____________ 3x +0 =12 divide by 3 x=4 substitute in 1st equation 4-y=10 subtract 4 from both sides -y=6 change sides by multiplying both sides by -1 y=-6 just a different way

OpenStudy (anonymous):

if you go to freemathhelp.com and scrolldown you will get an answer for any equations